# Box Confidence Depth: simulation-based inference with hyper-rectangles

### v2

@@ 3 Box-Confidence Depth
Assume that it is possible to generate data from the parametric model $p(y|\theta)$, with $\theta\in\Theta\subseteq\mathbb{R}^{p}$. Let $\theta_{0}$} be the true value of $\theta$. We assume that the model is correctly specified, so that $p(y|\theta_{0})$} corresponds to the true data generating process. Let $\pi(\theta)$ be a proposal distribution for the unknown parameter. Here, the proposal distribution is assumed to be uniform in a subset of the parameter space $\Theta^{b}\subset\Theta$. This boundedness of $\Theta^{b}$ is technically necessary to ensure computational feasibility, and guidelines to choose $\Theta^{b}$ are discussed below.
@@ 3.1 Scalar-scalar case
$\sup_{\theta\in\Theta^{b}}F_{t}(t^{\text{obs}}|\theta)>1-k\text{ and }\inf_{\theta\in\Theta^{b}}F_{t}(t^{\text{obs}}|\theta)<k,$
where $F_{t}(t^{\text{obs}}|\theta)$ is the cumulative distribution function of $t(y)$, computed at the value of the observed summary statistic $t^{\text{obs}}$ and with $k$ of the same order of the machine tolerance.
@@
Lemma 3.3.
Let $\hat{\theta}$ be the maximizer of $\mathcal{CD}^{\text{box}}(\theta)$. Then, under Assumption 1 $\hat{\theta}$ is median unbiased, i.e.
@@
Proof.
By definition $\hat{\theta}=\underset{\theta}{\arg\max}\>F_{t}(t^{\text{obs}}|\theta)[1-F_{t}(t^{\text{obs}}|\theta)].$ Since $F_{t}(t^{\text{obs}}|\theta)\in[0,1]$, the expression is maximum when the function x(1x)x(1-x) is maximum with $x$ in $[0,1]$, which is $F_{t}(t^{\text{obs}}|\hat{\theta})=0.5$. Then, applying $F_{t}$ to both sides of the inequality of Equation (1) it follows that $Pr_{\theta_{0}}(F_{t}(t^{\text{obs}}|\hat{\theta})<F_{t}(t^{\text{obs}}|\theta_{0}))=$ $Pr_{\theta_{0}}(0.5<F_{t}(t^{\text{obs}}|\theta_{0}))=1/2.$
@@ 3.5 Efficiency and optimality
where $\Theta_{1}$ denotes the parameter values under the alternative hypothesis. This inequality may be strict for some $\theta$, indicating that the CD-based test can achieve strictly higher power than the LRT in these cases. In other words, under the stated MLR and aggregation conditions, the CD-based test is uniformly at least as powerful as the LRT and may outperform it for certain alternatives.
@@ 3.6 High dimensional hyper rectangles
However, as stated in Section 3, a crucial property of distance and divergence functions typically used in ABC is the identity of indiscernible. This condition is not fulfilled by the type of discrepancy associated to the acceptance criterion of Box-CD. This would only occur in a degenerate case where the set $\mathcal{B}^{*}_{t}$ is a singleton with probability one, a scenario not endorsed by our assumptions. The failure to meet this property carries significant implications. In particular, it allows the method to accept parameter values even when the simulated summary statistics are far from the observed ones, as long as they fall within a coarse acceptance region. This does not rule out the problem of the curse of dimensionality. But we cannot study this issue straight under the lens of the distance-based approaches. To illustrate this, consider this simplified scenario. Let $t^{\text{obs}}\in\mathbb{R}^{d}$ be a fixed observed summary statistic, from the model $\mathcal{N}_{d}(0,I_{d})$ and let us assume that $t\sim\mathcal{N}_{d}(0,V)$ is a simulated statistic from the prior predictive distribution, where $V=v\cdot I_{d}$ is diagonal. In ABC, the common acceptance criterion is based on the fixed-radius ball
@@
$\mathbb{P}(\|t-t^{\text{obs}}\|\leq\epsilon)=\mathbb{P}((v+1)\cdot\chi^{2}_{d}\leq\epsilon^{2})=F_{\chi^{2}_{d}}\left(\frac{\epsilon^{2}}{v+1}\right),$
@@
$\mathbb{P}(\|t-t^{\textbf{obs}}\|\leq\epsilon)\approx p(t^{\text{obs}})\cdot Vol(\epsilon)=p(t^{\text{obs}})\cdot\frac{\pi^{d/2}}{\Gamma(d/2+1)}\cdot\epsilon^{d},$
@@
$\mathbb{P}\big(\|t_{1}-t_{2}\|\leq\epsilon\big)\propto\frac{1}{\sqrt{d}}\left(\epsilon\frac{\sqrt{2\pi e}}{\sqrt{d}}\right)^{d}<\epsilon^{d},$
which decays super-exponentially for the presence of $(1/\sqrt{d})^{d}$. Alternatively, consider the acceptance region defined by a data-adaptive axis-aligned hyperrectangle:
@@
$\mathbb{P}(t^{\text{obs}}\in\mathcal{B}(t^{(1)},t^{(2)}))\propto\frac{1}{\prod_{j=1}^{d}(1-x_{j})x_{j}}\leq\left(\frac{1}{4}\right)^{d},$
@@
The computational cost increases by $S/2$ due to the larger number of model simulations; however, the number of accepted values may grow faster than this rate, allowing for more accurate estimation of the target function—particularly in the tails—under a fixed computational budget. In particular, increasing $S$ offers a natural opportunity to exploit parallel computation: using $S$ parallel processors can in principle increase the probability of accepting beyond a factor of $S$. This contrasts with standard ABC rejection techniques, which typically cannot exploit multiple proposals as effectively. In ABC rejection, imposing a joint distance condition on multiple proposals alters the likelihood and decreases the acceptance probability; even when multiple model proposals are generated in parallel, and conditions are imposed independently, the number of accepted samples can only scale linearly with the number of proposals. In Example 4.3 we empirically examine the effect of choosing $S>2$, with particular attention to scenarios where the dimension of summary statistics increases.
@@ 4 Examples
where $h>0$ is the bandwidth parameter, internally chosen via cross-validation, and $K(\mathbf{u})=\prod_{j=1}^{d}k(u_{j})$, with $k(u_{j})$ representing the univariate Gaussian kernel function independently applied across each dimension $j$. We employed a 1-Nearest neighbor method to assess whether $\theta_{0}$ was included in the confidence regions, by predicting the value of $\mathcal{CD}^{\text{box}}(\theta_{0})$.
@@ 4.1 Logistic regression
Consider a logistic regression model for a sample of size $n=20$ with $p=3$ predictors and corresponding coefficients equal to $\beta_{0}=(-0.25,0,0.25)$. The summary statistics employed comprise the model’s sufficient statistics $t=X^{\top}y$, of dimension $d=p$. As a proposal, we use $\pi(\beta)=\text{Uniform}[-6,6]^{p}$. The empirical coverage level of confidence sets are closer to their nominal value than those obtained via the Likelihood Ratio test (Table 1).
@@ 4.2 Multivariate $t$ distribution
The true data generating parameter was set to $\mu_{0}=(0,-0.5,0.5)$. As a proposal we use $\text{Uniform}[-5,5]^{3}$ and as summary statistics the empirical medians of the components. The number of pseudo-samples generated for each parameter proposal was $S=2$. The results in Table 1 show that the method guarantees nominal coverage for confidence regions.
@@ 4.4 Ricker’s Model
We conduct two experiments: first, we assume that only the log-growth rate is unknown and consider as summary statistics the median of counts and the quantiles of level 0.25, 0.75. For the second experiment, both the parameters $log(r)$ and $\sigma^{2}$ were considered unknown, and we used as the set of summary statistics the whole time series minus the first observation, thus of length $d=19$. The number of pseudo-samples for each proposals were $S=2$ and $S=10$ in the two experiments, respectively. The empirical coverages of the confidence sets are conformal with the nominal (Table 1). Two examples of confidence regions obtained for two independent draws from the model with parameters $log(r)=2$ and $\sigma^{2}=2$ are reported in Figure 5.
@@ 5 Discussion
The first author acknowledges funding from the European Union under the ERC grant project number 864863.

### v3

@@ 3 Box-Confidence Depth
Assume that it is possible to generate data from the parametric model $p(y|\theta)$, with $\theta\in\Theta\subseteq\mathbb{R}^{p}$. Let $\theta_{0}$} be the true value of $\theta$. We assume that the model is correctly specified, so that $p(y|\theta_{0})$} corresponds to the true data generating process. Let $\pi(\theta)$ be a proposal distribution for the unknown parameter. Here, the proposal distribution is assumed to be uniform in a bounded subset of the parameter space $\Theta^{b}\subset\Theta$. The boundedness of $\Theta^{b}$, even when the parameter space is naturally unbounded, is technically necessary to ensure computational feasibility, and guidelines to choose $\Theta^{b}$ without compromising the validity of the inferential procedure are discussed below.
@@ 3.1 Scalar-scalar case
$\sup_{\theta\in\Theta^{b}}F_{t}(t^{\text{obs}}|\theta)>1-b\text{ and }\inf_{\theta\in\Theta^{b}}F_{t}(t^{\text{obs}}|\theta)<b,$
where $F_{t}(t^{\text{obs}}|\theta)$ is the cumulative distribution function of $t(y)$, computed at the value of the observed summary statistic $t^{\text{obs}}$ and with $b$ of the same order of the machine tolerance.
The set $\Theta_{b}$ serves a purely technical purpose: it guarantees that all parameter values with acceptance probability exceeding machine precision are included.
@@
Remark 1.
Let $\hat{\theta}$ be the maximizer of $\mathcal{CD}^{\text{box}}(\theta)$. Then, under Assumption 1, $\hat{\theta}$ is median unbiased, i.e.
@@
Indeed, by definition $\hat{\theta}=\underset{\theta}{\arg\max}\>F_{t}(t^{\text{obs}}|\theta)[1-F_{t}(t^{\text{obs}}|\theta)]$ and, since $F_{t}(t^{\text{obs}}|\theta)\in[0,1]$, the expression is maximum when the function x(1x)x(1-x) is maximum with $x$ in $[0,1]$, which is $F_{t}(t^{\text{obs}}|\hat{\theta})=0.5$. Then, applying $F_{t}$ to both sides of the inequality of Equation (1) it follows that $Pr_{\theta_{0}}(F_{t}(t^{\text{obs}}|\hat{\theta})<F_{t}(t^{\text{obs}}|\theta_{0}))=$ $Pr_{\theta_{0}}(0.5<F_{t}(t^{\text{obs}}|\theta_{0}))=1/2.$
@@ 3.5 Efficiency and optimality
where $\Theta_{1}$ denotes the parameter values under the alternative hypothesis. This inequality may be strict for some $\theta$, indicating that the CD-based test can achieve strictly higher power than the LRT in these cases. In other words, under the stated MLR and aggregating information from multidimensional statistics, the CD-based test is uniformly at least as powerful as the LRT and may outperform it for certain alternatives.
@@
Remark 6.
The conclusion of Theorem 3.7 should not be interpreted as a claim of classical optimality. In particular, the CD-based test is not asserted to be uniformly or locally most powerful in the Neyman-Pearson sense, where such optimality relies on monotone LR and sufficiency assumptions. Rather, the result establishes a containment relationship between rejection regions, implying power equivalence or dominance with respect to the LRT more in general. Importantly, the classical optimality of the LRT does not preclude the existence of alternative procedures that can match its local power, nor does it rule out improved performance in nonregular settings.
@@ 3.6 High dimensional hyper rectangles
However, as stated in Section 3, a crucial property of distance and divergence functions typically used in ABC is the identity of indiscernible. This condition is not fulfilled by the type of discrepancy associated to the acceptance criterion of Box-CD. This would only occur in a degenerate case where the set $\mathcal{B}^{*}_{t}$ is a singleton with probability one, a scenario not endorsed by our assumptions. The failure to meet this property carries significant implications. In particular, it allows the method to accept parameter values even when the simulated summary statistics are far from the observed ones, as long as they fall within a coarse acceptance region. This does not rule out the problem of the curse of dimensionality. But we cannot study this issue straight under the lens of the distance-based approaches.
3.6.1 ABC for increasing dimension
To illustrate this, consider this simplified scenario. Let $t^{\text{obs}}\in\mathbb{R}^{d}$ be a fixed observed summary statistic, from the model $\mathcal{N}_{d}(0,I_{d})$ and let us assume that $t\sim\mathcal{N}_{d}(0,V)$ is a simulated statistic from the prior predictive distribution, where $V=v\cdot I_{d}$ is diagonal. In ABC, the common acceptance criterion is based on the fixed-radius ball
@@ 3.6.1 ABC for increasing dimension
${Pr}(\|t-t^{\text{obs}}\|\leq\epsilon)={Pr}((v+1)\cdot\chi^{2}_{d}\leq\epsilon^{2})=F_{\chi^{2}_{d}}\left(\frac{\epsilon^{2}}{v+1}\right),$
@@
${Pr}(\|t-t^{\textbf{obs}}\|\leq\epsilon)\approx p(t^{\text{obs}})\cdot Vol(\epsilon)=p(t^{\text{obs}})\cdot\frac{\pi^{d/2}}{\Gamma(d/2+1)}\cdot\epsilon^{d},$
@@
${Pr}\big(\|t_{1}-t_{2}\|\leq\epsilon\big)\propto\frac{1}{\sqrt{d}}\left(\epsilon\frac{\sqrt{2\pi e}}{\sqrt{d}}\right)^{d}<\epsilon^{d},$
which decays super-exponentially for the presence of $(1/\sqrt{d})^{d}$.
3.6.2 Box-CD for increasing dimension
Alternatively, consider the acceptance region defined by a data-adaptive axis-aligned hyperrectangle:
@@ 3.6.2 Box-CD for increasing dimension
${Pr}(t^{\text{obs}}\in\mathcal{B}(t^{(1)},t^{(2)}))\propto\frac{1}{\prod_{j=1}^{d}(1-x_{j})x_{j}}\leq\left(\frac{1}{4}\right)^{d},$
@@ 3.6.3 Box-CD Extension with $S$ pseudo-samples
The computational cost of a single model simulation, and hence of one acceptance–rejection attempt, is $O(Sd)$ instead of $O(2d)$. However, the number of accepted samples may increase at a rate exceeding the ratio $S/2$, which can lead to more accurate estimation of the target function—especially in the tails—under a fixed computational budget. Moreover, increasing $S$ naturally lends itself to parallel computation: with $S$ parallel processors, the acceptance probability can grow by more than a factor of $S$ in practice. Finally, note that in general the cost of simulating from the model is typically higher than that of computing the statistics, thus $S$ dominates the cost for each proposal from the model.
In Example 4.3 we empirically examine the effect of choosing $S>2$, with particular attention to scenarios where the dimension of summary statistics increases.
@@ 4 Examples
where $h>0$ is the bandwidth parameter, internally chosen via cross-validation, and $K(\mathbf{u})=\prod_{j=1}^{d}k(u_{j})$, with $k(u_{j})$ representing the univariate Gaussian kernel function independently applied across each dimension $j$. We employed a 1-Nearest neighbor method to assess whether $\theta_{0}$ was included in the confidence regions, by predicting the value of $\mathcal{CD}^{\text{box}}(\theta_{0})$. The reported computing times in each example correspond to computations performed on a 3.49 GHz processor and parallelizing on 7 cores.
@@ 4.1 Logistic regression
Consider a logistic regression model for a sample of size $n=20$ with $p=3$ predictors and corresponding coefficients equal to $\beta_{0}=(-0.25,0,0.25)$. The summary statistics employed comprise the model’s sufficient statistics $t=X^{\top}y$, of dimension $d=p$. As a proposal, we use $\pi(\beta)=\text{Uniform}[-6,6]^{p}$. The empirical coverage level of confidence sets are closer to their nominal value than those obtained via the Likelihood Ratio test (Table 1). The computing time for $R=10000$ replications is 1.80 seconds.
@@ 4.2 Multivariate $t$ distribution
The true data generating parameter was set to $\mu_{0}=(0,-0.5,0.5)$. As a proposal we use $\text{Uniform}[-5,5]^{3}$ and as summary statistics the empirical medians of the components. The number of pseudo-samples generated for each parameter proposal was $S=2$. The results in Table 1 show that the method guarantees nominal coverage for confidence regions. The computing time in this example for $R=10000$ replications is of 2.17 seconds.
@@ 4.4 Ricker’s Model
We conduct two experiments: first, we assume that only the log-growth rate is unknown and consider as summary statistics the median of counts and the quantiles of level 0.25, 0.75. For the second experiment, both the parameters $log(r)$ and $\sigma^{2}$ were considered unknown, and we used as the set of summary statistics the whole time series minus the first observation, thus of length $d=19$. The number of pseudo-samples for each proposals were $S=2$ and $S=10$ in the two experiments, respectively. The empirical coverages of the confidence sets are conformal with the nominal (Table 1). Two examples of confidence regions obtained for two independent draws from the model with parameters $log(r)=2$ and $\sigma^{2}=2$ are reported in Figure 5. The computing time for $R=10000$ replications in the forst experiment is 2.28 seconds, for the second experiment 4.52 seconds.
@@ 5 Discussion
The first author acknowledges funding from the European Union under the ERC grant project number 864863 and from the Severo Ochoa Programme for Centres of Excellence in R&D (Barcelona School of Economics CEX2024-001476-S), funded by MCIN/AEI/10.13039/501100011033.